OpenStudy (anonymous):

3x - 4y = 7 3x + 2y = -5 When the second equation is subtracted from the first, the resulting equation is _____. -6y = -12 -6y = 12 -2y = -12 2y = 12

OpenStudy (anonymous):

got c

OpenStudy (anonymous):

Its -6y= -12

OpenStudy (anonymous):

\[3x-4y=7\] - \[3x + 2y = -5\] Remember everything has an opposite charge if subtracting \[-3x - 2y = -(-5)\] So the problem is looks more like this \[3x-4y=7\] + \[-3x - 2y = 5\] And you get \[-6y= -12\]